# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v5

@@ 5 Discussions
In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server. The server returns $N_{k}^{p}$ retrieved public instructions to the client $c_{k}$. FedDCA follows standard FedIT procedures, efficiently exchanging LoRA parameters $\Delta\phi_{k}$ between clients and the server.
@@
Comparing FedDCA with other FedIT methods Zhang et al. (2024b); Ye et al. (2024b), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption Acar et al. (2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
6 Experiments
To demonstrate the effectiveness of FedDCA and its variant FedDCA, we conduct extensive experiments across various domains and with several baselines. For additional results and analysis, please refer to Appendix C.
6.1 Experimental Setup
Alpaca CodeAlpaca HumanEval MedAlpaca MMLU-Med
$N_{sample}$ 52,002 20,022 164 33,955 1,089
Metric - - Pass@1 - Acc
FinGPT FPB FiQA TFNS MathInstruct GSM8K
$N_{sample}$ 76,772 152 35 299 224,567 1,319
Metric - Acc Acc Acc - Exact Match
@@ 6.1 Experimental Setup
Table 1: Dataset information of each domain. The test set is marked by $\ddagger$. The public dataset is composed of these five train sets. We denote HumanEval and MMLU-Med as H-Eval and M-Med for short.
Dataset and Evaluation Metrics.
As shown in Table 1, to evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical. We set the number of clients to 10 by default, while each client has 100 local instructions and obtains 5k augmented public instructions from the server.
Baselines.
We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg McMahan et al. (2017), FedProx Li et al. (2020), SCAFFOLD Karimireddy et al. (2020) and FedAvgM Hsu et al. (2019). 2) Augmented methods, including Random Sampling, Direct Retrieval, LESS Xia et al. (2024), and Self-Instruct Wang et al. (2022). Specifically, Random Sampling randomly samples $N_{k}^{p}$ public data for each client, while Direct Retrieval is a simplified version of FedDCA that directly retrieves public data based on each client’s cluster center set without greedy client center selection, not considering the cross-client domain coverage.
6.2 Main Results
Performance.
Table 2 shows the performance and the corresponding domain coverage of FedDCA and other baselines on four domains. FedDCA outperforms the other nine baselines in all domains, with a substantial improvement from at least 0.84% to the maximum of 29.19% over other baselines. In particular, FedDCA+FedAvg and FedDCA+FedProx perform better across four domains. We see that although FedDCA has a performance drop than FedDCA, it still outperforms other baselines. We can observe that FedIT methods perform well in the code domain, even better than most augmented methods. This could be attributed to two reasons: 1) The code domain follows a certain paradigm. 2) With limited local data, unaugmented methods learn the same data more times given the same epoch and batch size, which may be an unfair comparison. Nonetheless, FedDCA still surpasses the best baseline in the code domain.
Domain Coverage.
FedDCA achieves the highest domain coverage across all domains, surpassing other baselines by 4.82% to 21.36% on average. Although FedDCA uses a smaller encoder on the client side for computational efficiency, it results in a slight 2.89% drop in average domain coverage but still outperforms the best baseline by 2.78% on average. These results demonstrate the effectiveness of both FedDCA and FedDCA in enhancing domain coverage.
Method H-Eval M-Med FPB FiQA TFNS GSM8K
Zero-shot 29.88/- 70.60/- 55.94/- 18.54/- 59.21/- 23.27/-
FedAvg McMahan et al. (2017) 39.03/0.8126 68.40/0.6990 58.25/0.8529 14.18/0.8529 66.62/0.8529 47.46/0.7871
FedProx Li et al. (2020) 37.20/- 69.10/- 56.51/- 14.90/- 66.45/- 47.15/-
SCAFFOLD Karimireddy et al. (2020) 37.80/- 70.20/- 62.71/- 15.27/- 66.49/- 49.27/-
FedAvgM Hsu et al. (2019) 32.32/- 64.70/- 68.14/- 29.27/- 70.32/- 46.85/-
Random Sampling 32.93/0.8512 71.30/0.7940 64.19/0.9196 13.09/0.9196 65.53/0.9196 47.38/0.8651
Direct Retrieval 34.14/0.9396 72.20/0.8830 66.31/0.9293 19.11/0.9293 67.62/0.9293 50.87/0.8967
LESS Xia et al. (2024) 28.04/0.8509 71.00/0.7737 60.56/0.8917 16.00/0.8917 61.14/0.8917 43.13/0.8352
Self-Instruct Wang et al. (2022) 32.92/0.8966 71.90/0.8586 59.73/0.9015 20.67/0.9015 66.54/0.9015 50.79/0.8811
FedDCA 34.75/0.9532 73.30/0.8972 67.10/0.9538 30.54/0.9538 71.01/0.9538 51.55/0.9096
FedDCA+FedAvg 36.58/0.9766 74.50/0.9348 67.24/0.9815 35.27/0.9815 73.32/0.9815 52.46/0.9320
FedDCA+FedProx 32.92/- 72.40/- 72.93/- 38.18/- 77.55/- 51.25/-
FedDCA+SCAFFOLD 39.87/- 73.20/- 72.68/- 33.09/- 75.50/- 50.26/-
FedDCA+FedAvgM 33.53/- 68.90/- 71.45/- 31.45/- 72.52/- 49.76/-
Table 2: Performance (%) and domain coverage of FedDCA, FedDCA and other nine baselines. Due to limited local data or inappropriate data augmentation strategy, occurs performance degradation in the Unaugmented and Random Sampling settings (underlined values). FedDCA outperforms all other baselines.
Plug-and-Play.
To demonstrate the plug-and-play capability of FedDCA with other augmentation methods, we first perform instruction sampling through different baselines to obtain augmented data (5k instructions per client) and then utilize FedDCA to conduct data selection (select 1k instructions per client) for FedDIT. As shown in Table 3, using FedDCA in combination with other augmentation methods still achieves comparable performance despite using only 20% of the sampled data. FedDCA efficiently selects the most representative samples based on domain coverage, enabling effective training with less data while maintaining performance.
Computation Analysis.
To show the computational efficiency of FedDCA, we compare these to FedDCA from two aspects: 1) Encoder model size. 2) Encoding time overhead. We test the encoders’ encoding time on the public dataset. As shown in Figure 6(a), FedDCA requires only about 30% encoded time overhead of FedDCA. Moreover, FedDCA achieves a substantial reduction in model size, which is less than 10% of FedDCA’s.
Privacy Analysis.
We evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks Carlini et al. (2021); Zhang et al. (2024a), which utilizes the autoregression nature of LLMXu et al. (2024). We analyze the domain inference attack in Section C.3.
We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1k to 5k public instructions. We set up 10 clients with full participation for 10 rounds. Specifically, we record the average ROUGE-L score Lin (2004) denoted as $\rho$ for client $c_{0}$ in each round. The prompt to extract instructions is in Section C.1. For each setting, we repeat memory extraction 100 times and report $\rho$. Specifically, denote the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$. The calculation is defined as $\rho=\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
@@ 6.2 Main Results
(b) $\rho$ per round. Dotted lines represent the base-data-only setting.

### v6

@@ A.1 Discussions
The communication overhead of FedDCA is mainly incurred in two phases: (1) In the domain instruction augmentation phase, each client uploads $\xi$ cluster centers (typically low-dimensional vectors, e.g., 1024-d) to the server, and the server returns $N_{k}^{p}$ retrieved public instructions to each client. (2) During model parameter synchronization, FedDCA follows standard federated learning (FedIT) procedures, requiring only the exchange of LoRA parameters $\Delta\phi_{k}$, which significantly reduces communication cost.
@@
Comparing FedDCA with other FedIT methods Zhang et al. (2024c); Ye et al. (2024c), the difference lies in the client center selection stage. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption Acar et al. (2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
Robustness to Outliers.
The k-means clustering, a component of our client-side processing, can be sensitive to outliers in the local data. If a client’s local dataset contains significant outliers, these could potentially skew the resulting cluster centers, and subsequently affect the server-side selection and retrieval process. To mitigate this, two main approaches could be considered. Firstly, more robust clustering algorithms from existing literature could be adopted by clients. For instance, Deshpande and Pratap, 2023 enhances k-means++’s robustness by capping the selection probability for new centers, based on the overall clustering cost and an allowed number of outliers. This prevents distant, potentially outlier, points from dominating the initialization process and improves clustering in noisy data. Secondly, clients could perform preliminary data filtering or cleaning steps before initiating the FedDCA process. This might involve simple heuristic-based filtering (e.g., keyword-based removal of irrelevant instructions) or leveraging lightweight LLM-based tagging to identify and exclude potential outliers Bernsohn et al. (2024); Biester et al. (2024); Zhang et al. (2024a). A comprehensive investigation into the impact of outliers and the optimal strategies for enhancing FedDCA’s robustness in such scenarios is a promising direction for future work.
A.2 Theoretical Foundation of FedDCA
FedDCA’s optimization objective is equivalent to the classic submodular facility location problem, ensuring our greedy algorithm achieves at least a 11/e1-1/e approximation ratio.
The Optimization Problem in FedDCA.
The overall optimization problem in our paper is formulated as Eq. 3. In our specific problem setting, the number of cluster centers ($\xi$) per client is considered fixed. Therefore, minimizing the objective function is equivalent to maximizing the domain coverage term $d(D^{d},\mathcal{P})$ through the selected client centers $\mathcal{P}$.
The Submodular Facility Location Problem.
In its abstract form, we are given: a set of “locations” or “customers” $\mathcal{U}$ that need to be serviced; a set of potential “facilities” $\mathcal{F}$ that can be opened; a benefit function $w(i,j)$ that quantifies the value of servicing customer $j\in\mathcal{U}$ with facility $i\in\mathcal{F}$. The objective is to select a subset of facilities $\mathcal{S}\subseteq\mathcal{F}$ of a given size $k$ to maximize the total benefit provided to all customers:
$g(\mathcal{S})=\sum_{j\in\mathcal{U}}\max_{i\in\mathcal{S}}w(i,j)$ (4)
Formal Equivalence.
The connection between our domain coverage problem and the facility location problem becomes clear when we map the components: 1) Customers ($\mathcal{U}$) $\rightarrow$ The set of in-domain data, $D^{d}$; 2) Facilities ($\mathcal{F}$) $\rightarrow$ The set of all candidate centers, $\mathcal{C}_{\text{all}}$; 3) Benefit ($w(i,j)$) $\rightarrow$ The cosine similarity between two centers, $\operatorname{sim}(p,c)$.
Candidate centers are obtained by clustering each client’s local data, whereas $D^{d}$ is the union of all public in-domain data points. This “non-diagonal” mapping still matches the standard facility-location structure because every data point is evaluated against the chosen center set, regardless of origin.
With this mapping, our domain coverage objective $d(D^{d},\mathcal{P})$ has the form of the classic facility location function $g(\mathcal{P})$ normalized by the size of the in-domain data, $|D^{d}|$. That is, $d(D^{d},\mathcal{P})=\frac{1}{|D^{d}|}g(\mathcal{P})$.
Method
Privacy
Preserving
API
Cost
Additional
Information
Performance
Degradation
Domain Coverage
Oriented
Computational Overhead of
Data Augmentation
FedAvg McMahan et al. (2017) -
Random Sampling -
LESS Xia et al. (2024) high
FewFedPIT Zhang et al. (2024c) high
KnowledgeSG Wang et al. (2024) high
Self-Instruct Wang et al. (2022) low
Direct Retrieval low
FedDCA (ours) low
Dataset Size Metric
Alpaca 52,002 -
MedAlpaca 33,955 -
MMLU-Med 1,089 Acc.
FinGPT 76,772 -
FPB 152 Acc.
FiQA 35 Acc.
TFNS 299 Acc.
MathInstruct 224,567 -
GSM8K 1,319 Exact Match
@@ A.2 Theoretical Foundation of FedDCA
Table 5: Dataset information of each domain. Public dataset is composed of these five train sets (Alpaca, MedAlpaca, FinGPT, MathInstruct, and GSM8K). Test sets are marked with $\dagger$ and used for evaluation.
Monotonicity and Submodularity.
The facility location function $g(\mathcal{P})$ is known to be both monotone and submodular. Our objective $d(D^{d},\mathcal{P})$ inherits these properties.
Monotonicity: This property is guaranteed because our benefit function, $\operatorname{sim}(p,c)$, is the cosine similarity from a sentence encoder, which is non-negative [^1] https://huggingface.co/BAAI/bge-large-en-v1.5 (we have plotted histograms of the similarity scores for each domain, which fall within the range about [0.2, 0.85]). When adding a new center to $\mathcal{P}$, the maximum similarity for any data point $c$ can only increase or stay the same. If a negative value ever arises, we can rectify the score via applying an affine shift, which preserves the order of similarities and therefore the monotonicity. Consequently, the total sum of maximum similarities, $g(\mathcal{P})$, cannot decrease.
Submodularity (Diminishing Returns): To formally show this, let’s define the marginal gain for a single data point $c\in D^{d}$ when adding a new center $x$ to a set of centers $S$ as:
$\Delta_{c}(S,x)=\max\bigl{(}0,\operatorname{sim}(x,c)-\max_{p\in S}\operatorname{sim}(p,c)\bigr{)}$ (5)
This formula calculates the new coverage a center $x$ provides to a point $c$, beyond what is already provided by the existing set $S$. Now, consider two sets of centers $A$ and $B$ such that $A\subseteq B$. For any data point $c$, the maximum similarity from set $B$ must be at least as large as from set $A$, so:
$\max_{p\in A}\operatorname{sim}(p,c)\leq\max_{p\in B}\operatorname{sim}(p,c)$ (6)
Consequently, the marginal gain from adding $x$ must be smaller for the larger set $B$. That is, for any $c$:
$\Delta_{c}(A,x)\geq\Delta_{c}(B,x)$ (7)
The total marginal gain of adding center $x$ to a set $S$ is:
$g(S\cup\{x\})-g(S)=\sum_{c\in D^{d}}\Delta_{c}(S,x)$ (8)
Since the inequality holds for each individual term in the sum, it must hold for the sum itself. This demonstrates the diminishing returns property:
$g(A\cup\{x\})-g(A)\geq g(B\cup\{x\})-g(B)\text{ for }A\subseteq B$ (9)
Overall, our problem of maximizing the domain coverage $d(D^{d},\mathcal{P})$ is equivalent to solving the facility location problem of maximizing $g(\mathcal{P})$, ensuring theoretical guarantees for FedDCA.
In addition, we have provided the time complexity analysis in Section˜A.1 and our empirical results show that the FedDCA algorithm converges quickly (about 1.1N passes). Furthermore, Section˜B.9 demonstrates that FedDCA achieves near-optimal performance in practice.
Appendix B Experiments
B.1 Train and Test Dataset Information
We evaluate our method on multiple domains including medical (MedAlpaca, MMLU-Med), financial (FinGPT, FPB, FiQA, TFNS), and mathematical reasoning (MathInstruct, GSM8K). Table 5 shows the dataset statistics for each domain. The public dataset used for retrieval consists of the training sets from these domains and the general instruction dataset (Alpaca) through concatenation and random shuffling, while the test sets (marked with $\dagger$) are used for evaluation.
B.2 Baselines
To highlight FedDCA’s advantages, Table 4 provides a comparative overview with existing baseline methods along six critical dimensions: Privacy Preserving, API Cost, Additional Information requirements, potential for Performance Degradation, orientation towards Domain Coverage, and the Computational Overhead of Data Augmentation. These dimensions are crucial for assessing the practical viability of federated instruction tuning strategies. We analyze each dimension in detail below:
Figure 5: Prompt used for memory extraction attack.
  • Privacy Preserving: Most evaluated methods, including FedDCA, inherently preserve client data privacy by design, as they do not share raw local data. Self-Instruct is a notable exception as it involves sending prompts containing potentially sensitive information to an external API. Our detailed privacy analysis for FedDCA can be found in Sections˜A.1 and B.5.
  • API Cost: FedDCA operates without requiring external API calls, thus incurring no associated costs. In contrast, Self-Instruct, by leveraging large models like GPT for instruction generation, entails API expenses.
  • Additional Information: FedDCA primarily relies on clients’ local training data for its operations. Some methods, such as LESS Xia et al. (2024), necessitate additional resources like a validation set for its gradient-based retrieval mechanism.
  • Performance Degradation Avoidance: A key objective is to enhance model performance or, at minimum, avoid degradation. As indicated in our main results (Table 1), FedDCA is designed to prevent performance drops. In contrast, methods like FedAvg (when unaugmented), Random Sampling, and even some generation-based approaches like FewFedPIT and KnowledgeSG (if generation quality is not optimal or well-aligned) can risk performance degradation.
  • Domain Coverage Oriented: FedDCA is uniquely engineered to explicitly optimize cross-client domain coverage in distributed environments. This strategic focus is a distinguishing feature largely absent in the other compared baselines, which typically do not have a direct mechanism for maximizing instruction diversity across the federation.
  • Computational Overhead of Data Augmentation: FedDCA maintains a low computational overhead for its data augmentation phase, comparable to methods like LESS (due to gradient computation and warm-up), FewFedPIT (local LLM generation), and KnowledgeSG (expert model generation on the server) inherently involve higher computational demands for data augmentation.
In conclusion, these comparisons underscore FedDCA’s well-rounded design. It effectively addresses key practical challenges by preserving privacy, avoiding API costs, minimizing reliance on additional information, preventing performance degradation, strategically optimizing domain coverage, and maintaining a low computational overhead for augmentation.
B.3 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is 5e55e-5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means Wu (2012), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions.
Figure 6: Impact of different retrieval amounts on FedDCA’s performance (%) and domain coverage. The dotted lines represent the domain coverge. Results demonstrate FedDCA’s robust applicability across different augmentation scales.
Specifically, for LESS, we uses LoRA with rank 128 and dropout 0.1 for warmup training on a random 5% subset of the public dataset for 4 epochs. We apply a learning rate peak of $2\times 10^{-5}$ with linear warm-up and cosine decay. Gradients are extracted from each epoch checkpoint and projected to a 1024-dimensional space using Johnson-Lindenstrauss random projections. Data selection is performed by computing cosine similarity between training and validation gradient features aggregated across epochs. The top $N_{k}^{p}$ scoring examples are selected for final client $c_{k}$’s instruction tuning. Additionally, for KnowledgeSG, we utilize AlpacaCare Zhang et al. (2023d), FinGPT Yang et al. (2023a) and WizardMath Luo et al. (2023) as the server-hosted expert model for each domain respectively. To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 4. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
B.4 Effect of Retrieval Number
@@
(b) The average ROUGE-L score per round. Dotted lines represent the base-data-only setting.
@@
Experimental Setup.
To assess the impact of the volume of augmented data, we vary the number of public instructions retrieved per client by FedDCA. Specifically, we test retrieval amounts of [100, 200, 400, 800, 1000] instructions per client, while keeping other experimental parameters consistent with the main setup described in Section˜6.1. The experiments are conducted across the three primary domains (medical, financial, and mathematical).
Results.
Figure˜6 shows FedDCA’s performance and domain coverage with varying retrieval amounts. Both generally increase with more retrieved instructions, indicating richer augmentation enhances learning. The improvement rates slow after about 400 instructions per client, likely due to increasing overlap with already covered semantic space or the model reaching its capacity to benefit from further diverse examples under the given training regime. Overall, performance and coverage continue to rise with retrieval volume.
Notably, domain coverage’s impact on performance differs by domain. For example, medical domain gains are modest, while financial domain gains are more aligned with coverage increases. This variance may stem from: 1) The base model’s initial domain proficiency, which influences its baseline performance and the marginal benefits of additional data. 2) The degree of similarity between the public in-domain data distribution and the test set distribution, where higher relevance yields more proportional performance benefits.
B.5 Privacy Analysis
Memory Extraction Attack.
We evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks Carlini et al. (2021); Zhang et al. (2024b), which utilizes the autoregression nature of LLMXu et al. (2024b).
We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 100 to 1k public instructions. We set up 10 clients with full participation for 10 rounds. Specifically, we record the average ROUGE-L score Lin (2004) for client $c_{0}$ in each round (following the setup in Zhang et al. (2024c)). As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 5, which is exactly the prefix of the Alpaca’s template.
For each setting, we repeat memory extraction 100 times and report the average ROUGE-L score. Specifically, denote the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$. The calculation of the average ROUGE-L score is defined as $\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.